OpenStudy (anonymous):

find the general solution for dy/dx-y=15 by integration

OpenStudy (lgbasallote):

this has been answered already why did you repost this??

OpenStudy (anonymous):

because i didn't get the correct answer

OpenStudy (anonymous):

i just looked at the solution given by @lalaly , why was it wrong? i don't see anything wrong with that solution...

OpenStudy (anonymous):

\[ \frac{dy}{dx}-y=15 \\ \frac{dy}{dx} =15 +y \\ \frac{dy}{y+15}= dx \\ \int \frac{dy}{y+15}= \int dx\\ \ln(y+15) = x + K \\ e^{\ln(y+15)} = e^{x + K }= e^K e^x= C e^x\\ y+15= Ce^x \\ y= Ce^x -15 \]

OpenStudy (anonymous):

And that's the same answer ^^...

OpenStudy (anonymous):

I did not look at the other answer. This is a routine method for separable differential equations.